OpenStudy (anonymous):

2+2= first to answer gets medal

OpenStudy (turingtest):

why, do you not know the answer?

hero (hero):

2 + 2 = 5 - 1

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4

hero (hero):

I answered first

OpenStudy (anonymous):

no hero the answer was 4 not 5-1

OpenStudy (anonymous):

yes.., haha

hero (hero):

Five minus one equals four

OpenStudy (anonymous):

yes Five minus one equals four

OpenStudy (anonymous):

@gerryliyana gave you a medal for the question: 2+2= first to answer gets medal @Hero

hero (hero):

That's nice :P